# PCHands: PCA-based Hand Pose Synergy Representation on Manipulators with N-DoF

### v1

@@ I Introduction
To overcome these limitations, we propose PCHands, a novel PCA-based method for extracting a common representation of hand postural synergies spanning human, dexterous robotic hands, and two-finger grippers. PCHands learns a variable-length postural synergy representation across a large set of very different manipulators by means of three components: the proposed Anchor Description Format (ADF), Conditional Variational Auto-Encoder (CVAE[14], and Principal Component Analysis (PCA[15]. In ADF, we first position a predefined set of anchors (also known as keypoints) on each manipulator. We then collect a dataset of anchor positions for all manipulators at random joint configurations. Next, we devise a procedure that iteratively (i) trains a CVAE on the anchors dataset to extract a common latent representation of the postural synergies, and (ii) aligns the end-effector of each manipulator with Iterative Closest Point (ICP). Moreover, we propose to further process the latent representation with PCA to extract a variable-length latent representation. This provides flexibility in dimensionality selection and allows for accommodating different levels of complexity in both manipulators and downstream tasks.
@@ IV-D Transfer to Real-World Setup
In this section, we examine the roll-out of the DAPG policies learned with PCHands in simulation in Sec. IV-C on a real robot. As depicted in Fig. 7, the considered platform includes a 7-DoF Franka-Panda robotic arm equipped with either a Robotiq-2f85 (2F in Table I) or a LEAP-hand (4F), and an external RGB-D camera (a RealSense L515). To fulfill the observation space of the policy, $T^{eef}$, $N$-pc and the object 6D pose $T^{obj}$ are needed. We get $T^{eef}$ and the $N$-pc manipulator configuration with proprioception, and use FoundationPose [39] as a 6D object tracker that provides the pose of the target object from RGB-D images. We then use the libfranka motion generator to map the output of our policy in the form of 6D end-effector poses to joint trajectories. Table I reports the success rate over 10 roll-outs achieved by the policies trained for 400 steps in Sec. IV-C on the three Relocate tasks, deployed zero-shot on the real robot without any sim-to-real adaptation.

### v2

@@ I Introduction
To overcome these limitations, we propose PCHands, a novel PCA-based method for extracting a common representation of hand postural synergies spanning human, dexterous robotic hands, and two-finger grippers. PCHands learns a variable-length postural synergy representation across a large set of very different manipulators by means of three components: the proposed Anchor Description Format (ADF), Conditional Variational Auto-Encoder (CVAE) [14], and Principal Component Analysis (PCA) [15]. In ADF, we first position a predefined set of anchors (also known as keypoints) on each manipulator. We then collect a dataset of anchor positions for all manipulators at random joint configurations. Next, we devise a procedure that iteratively (i) trains a CVAE on the anchors dataset to extract a common latent representation of the postural synergies, and (ii) aligns the end-effector of each manipulator with Iterative Closest Point (ICP). Moreover, we propose to further process the latent representation with PCA to extract a variable-length latent representation. This provides flexibility in dimensionality selection and allows for accommodating different levels of complexity in both manipulators and downstream tasks.
@@ IV-D Transfer to Real-World Setup
In this section, we examine the roll-out of the DAPG policies learned with PCHands in simulation in Sec. IV-C on a real robot. As depicted in Fig. 7, the considered platform includes a 7-DoF Franka-Panda robotic arm equipped with either a Robotiq-2f85 (2F in Table I) or a LEAP-hand (4F), and an external RGB-D camera (a RealSense L515). To fulfill the observation space of the policy, $T^{eef}$, NN-pc and the object 6D pose $T^{obj}$ are needed. We get $T^{eef}$ and the NN-pc manipulator configuration with proprioception, and use FoundationPose [39] as a 6D object tracker that provides the pose of the target object from RGB-D images. We then use the libfranka motion generator to map the output of our policy in the form of 6D end-effector poses to joint trajectories. Table I reports the success rate over 10 roll-outs achieved by the policies trained for 400 steps in Sec. IV-C on the three Relocate tasks, deployed zero-shot on the real robot without any sim-to-real adaptation.